# Mitigating Backdoor Attacks using Activation-Guided Model Editing

### v1

@@ 1 Introduction
Gu et al. [15] proposed BadNets, the first backdoor attack. The attack overlays a small subset of training samples with a square of fixed size and position, and it changes the labels to a target class, thus poisoning the samples. During training, the victim model learns to associate the trigger pattern with the target class, which creates a hidden backdoor reactive to the trigger. During inference, the model behaves as usual on clean data. Still, when a malicious actor forwards a sample with a specific trigger, the backdoor in the neural network is activated, leading to model misbehavior, such as misclassification. A survey from Microsoft stated that data poisoning is one of the top attacks on machine learning systems [27].
@@ 2.2 Backdoor Defenses
With the development of backdoor attacks, researchers have proposed various backdoor defenses as countermeasure [32]. Most defense methods should instead be considered mitigation methods because, in most cases, they cannot entirely erase the influence of the attacks. One such mitigation is data pre-processing prior to model inference, which aims to perturb the trigger to not activate the backdoor [38, 8]. Another form of defense, typically used to aid other defense methods, is trigger synthesis [51]. With reverse engineering, trigger synthesis approximates the trigger, which can be used for trigger-guided defense or to retrieve the target class. Model diagnosis [25, 55] is another type of defense to detect a backdoor and prevent model deployment. Moreover, poison suppression defenses modify the training process to be robust against backdoor creation [10, 20]. Other methods utilize sample filtering to detect trigger images and remove them from the training set or decline them during model inference [11, 50, 6]. Some defense methods aim to remove the backdoor from an infected model by directly modifying the model  [38, 35]. In this work, we propose a method that directly edits model weights to erase the backdoor in a poisoned model. The editing uses the activation of clean or poisoned examples as a guide.
@@ 3 Methodology
Given $f_{\theta}$ without having access to the training dataset $D$, we propose an activation-guided model editing approach to unlearn $D_{B}$ under two assumptions: (1) we have Backdoor Knowledge (BDK), and (2) we do not have it ($\neg$¬BDK). For both assumptions, we split the total weights of $f_{\theta}$ into two halves and add those layers corresponding to the weights of the second half to a layer list $L$, for which we want to edit the weights. We target those later layers because they have the highest proximity to the classification output. We do not want to edit the early layers associated with general low-level feature extraction [13]. First, we prepare an unlearning dataset $D_{U}$ with the same distribution as the training dataset $D$. However, $D_{U}$ is not used in training $f_{\theta}$. Our empirical experiments suggest that $D_{U}$ can be as small as four samples.
@@ 3.1 Assumption 1 - BDK
Extract model activation $A$ for the whole dataset $D_{U}$. Therefore, iterate batch-wise over data $X$ in $D_{U}$ infected with trigger $\delta$ and forward it through model $f_{\theta}$ given $f_{\theta}(X+\delta)$. Capture and average the activations across all batches.
@@
Then, iterate over each layer $l$ in the layer list $L$ of the layers target for editing. For each layer $l$, check if the multiplication of weights $\theta_{l}$ and activations $A_{l}$ would result in a matrix-multiplication-error caused by a shape mismatch. If this is the case, utilize adaptive average pooling to expand the shape of $A_{l}$ to correspond to the shape of $\theta_{l}$.
@@
In Equation 4, we rescale the activation with values 0.5 and 1.0 for $\gamma$ and $\lambda$, which shifts the mean of the activation to 1.0. By doing this, we aim to preserve the utility of the model. We hypothesize that this leads to minimal change in the inherent mean of the weights when multiplying the rescaled activations with the weights. We experimentally confirmed the performance with those values in the supplementary material.
After model editing, we observe that the model tends to unlearn the whole class instead of unlearning backdoored samples only. To address this issue, we introduce an optional repair phase to restore some model utility if we have unlearned more than intended. Specifically, we fine-tune the model one time on $D_{U}$ and another time on $D_{U}$ poisoned with $\delta$ for one epoch.
@@ 3.2 Assumption 2 - $\neg$¬BDK
Table 6: Efficiency of proposed unlearning compared with state-of-the-art methods. We experimented with 50%(5000), 5%(500), 0.5%(50), and 0.05%(5) of unseen CIFAR10 data for unlearning. Table displays only sample runs with highest and second-highest scores. Full table with all sample runs is displayed in supplementary material. Best results are highlighted in bold.
@@ 4.1 Setup
Baselines. We considered four backdoor unlearning methods for comparison, two of which require BDK: (1) fine-tuning, which penalizes a high difference between activations of clean and poisoned data (actFT) [43], and (2) BaEraser, which uses gradient ascent for unlearning  [36]. The other two methods work with $\neg$¬BDK: (3) fine-tuning on clean $D_{U}$ in the same way as the initial training (basicFT), and (4) Neural Attention Distillation (NAD), a knowledge distillation approach where the basicFT model, acting as the teacher model, only passes on its ability to clean data [31].
@@
Score $=$ $(1-\frac{\textrm{ASR}^{U}}{\textrm{ASR}^{V}})\cdot\frac{\textrm{ACC}^{U}}{\textrm{ACC}^{R}}.$ (6)
Base Configuration. We use specific base configurations if not stated otherwise for an experiment. Experiments were performed on the CIFAR10 dataset and ResNet18 as the victim model. The training dataset $D_{T}$ was infected with a poisoning rate $\rho$ of 5%, with the trigger displayed in Figure 2(a). The backdoor target class $y_{t}$ is two, representing birds. For the activation scale and shift hyperparameters $\gamma$ and $\lambda$, we use values 0.5 and 1.0, respectively. The ablation study for obtaining those values is discussed in the supplementary material. 5000 samples are usually available for unlearning dataset $D_{U}$, but our method only uses 512 by default.
@@ 4.2 Results
Tables 3 and 4 summarize the performance of the proposed unlearning method with the different baseline methods in terms of score (see Section 4.1). The score metric measured the forgetting and utility quality after unlearning. For backdoor verification, the $alpha$ value was multiplied by three up to a maximum of 100%. For actFT to be effective, we multiplied the $alpha$ value for unlearning by the same magnitude. Our method outperformed the previous methods in terms of score with or without BDK for most triggers. The basic fine-tuning with $\neg$¬BDK achieved a higher score for the Narcissus trigger.
@@ 5 Discussion
We experimented with our algorithm and found specific activation-manipulating formulas that gave us the best unlearning performance for model editing. In Equation 1, negating poisoned activation with BDK and clean activation with $\neg$¬BDK worked the best. With BDK, we negate the activation of the trigger-infected data we want to forget. Previously, Ilharco et al. [22] came to the same conclusion that using the negative for information leads to unlearning.
@@ 0.B.3 Effect of Activation Mean and Standard Deviation
Figure 4: Performance of our method when using different values for hyperparameters $\gamma$ and $\lambda$. Compared with or without change in BN MA parameters.

### v2

@@ 1 Introduction
Gu et al. [15] proposed BadNets, the first backdoor attack. BadNets overlays a small subset of training samples with a square of fixed size and position, and it changes the labels to a target class, thus poisoning the samples. During training, the victim model learns to associate the trigger pattern with the target class, creating a hidden backdoor reactive to the trigger. During inference, the model behaves as usual on clean data. Still, when a malicious actor forwards a sample with a specific trigger, the backdoor in the neural network is activated, leading to model misbehavior, such as misclassification. A survey from Microsoft stated that data poisoning is one of the top attacks on machine learning systems [27].
@@ 2.2 Backdoor Defenses
With the development of backdoor attacks, researchers have proposed various backdoor defenses as countermeasure [32]. Most defense methods should instead be considered mitigation methods because, in most cases, they cannot entirely erase the influence of the attacks. One such mitigation is data pre-processing prior to model inference, which aims to perturb the trigger to not activate the backdoor [38, 8]. Another form of defense, typically used to aid other defense methods, is trigger synthesis [51]. With reverse engineering, trigger synthesis approximates the trigger, which can be used for trigger-guided defense or to retrieve the target class. Model diagnosis [25, 55] is another type of defense to detect a backdoor and prevent model deployment. Moreover, poison suppression defenses modify the training process to be robust against backdoor creation [10, 20]. Other methods utilize sample filtering to detect trigger images and remove them from the training set or decline them during model inference [11, 50, 6]. Some defense methods aim to remove the backdoor from an infected model by directly modifying the model  [38, 35]. In this work, we propose a method that directly edits model weights to erase the backdoor in a poisoned model. The editing uses the activation of clean or poisoned samples as a guide.
@@ 3 Methodology
Given $f_{\theta}$ without having access to the training dataset $D$, we propose an activation-guided model editing approach to unlearn $D_{B}$ under two assumptions: (1) we have Backdoor Knowledge (BDK), and (2) we do not have it ($\neg$¬BDK). For both assumptions, we split the total weights of $f_{\theta}$ into two halves and add those layers corresponding to the weights of the second half to a layer list $L$, for which we want to edit the weights. We target those later layers because they have the highest proximity to the classification output. We do not want to edit the early layers associated with general low-level feature extraction [13]. The authors of other backdoor mitigation approaches also suggest that focusing the unlearning on the later layers improves performance [35, 51]. First, we prepare an unlearning dataset $D_{U}$ with the same distribution as the training dataset $D$. However, $D_{U}$ is not used in training $f_{\theta}$. Our empirical experiments suggest that $D_{U}$ can be as small as four samples.
@@ 3.1 Assumption 1 - BDK
Extract model activation $A$ for the whole dataset $D_{U}$. Therefore, iterate batch-wise over data $X$ in $D_{U}$ infected with trigger $\delta$ and forward it through model $f_{\theta}$ given $f_{\theta}(X+\delta)$. Capture and average the activations across all batches for each layer $l$ in layer list $L$.
@@
Then, iterate over $l$ in list $L$ of the layers targeted for editing. For each layer $l$, check if the later multiplication of weights $\theta_{l}$ and activation $A_{l}$ would result in a matrix-multiplication-error caused by shape mismatch. If this is the case, use adaptive average pooling to adjust the shape of $A_{l}$ to match the shape of $\theta_{l}$.
@@
In Equation 4, we scale and shift the activation with values 0.5 and 1.0 for $\gamma$ and $\lambda$ hyperparameters, which changes the activation mean to 1.0. With this, we aim to preserve the utility of the model during unlearning and mitigate the internal covariate shift, especially after significant distribution changes by Equation 1. Choosing those values should lead to minimal change in the inherent mean of the weights when multiplying with the shifted activation. We experimentally confirmed the performance with those values in the supplementary material.
After model editing, we observe that the model tends to unlearn the whole class instead of unlearning backdoored samples only. To address this issue, we introduce an optional repair phase to restore some model utility if we have unlearned more than intended. Specifically, we fine-tune the model on the samples in $D_{U}$ and on the same samples poisoned with $\delta$ for one epoch each, using the correct ground truth as the target label in both cases.
@@ 3.2 Assumption 2 - $\neg$¬BDK
Table 6: Efficiency of proposed unlearning compared with state-of-the-art methods. Experimented with 50%(5000), 5%(500), 0.5%(50), and 0.05%(5) of unseen CIFAR10 data for unlearning. Table displays only sample runs with highest and second-highest scores. Full table is displayed in supplementary material. Best results are highlighted in bold.
@@ 4.1 Setup
Baselines. We considered four suitable backdoor unlearning methods for comparison, two of which require BDK: (1) fine-tuning, which penalizes a high difference between activations of clean and poisoned data (actFT) [43], and (2) BaEraser, which uses gradient ascent for unlearning  [36]. The other two methods work with $\neg$¬BDK: (3) fine-tuning on clean $D_{U}$ in the same way as the initial training (basicFT), and (4) Neural Attention Distillation (NAD), a knowledge distillation approach where the basicFT model, acting as the teacher model, only passes on its ability to clean data [31].
@@
$\textrm{Score}=(1-\frac{\textrm{ASR}^{U}}{\textrm{ASR}^{V}})\cdot\frac{\textrm{ACC}^{U}}{\textrm{ACC}^{R}}.$ (6)
Base Configuration. We used specific base configurations if not stated otherwise for an experiment. Experiments were performed on the CIFAR10 dataset and ResNet18 as the victim model. The training dataset $D_{T}$ was infected with a poisoning rate $\rho$ of 5%, with the trigger displayed in Figure 2(a). The backdoor target class $y_{t}$ was two, representing birds. Unlearn dataset $D_{U}$ consisted of 5000 samples, but our method only used 512 by default.
@@ 4.2 Results
Tables 3 and 4 summarize the performance of the proposed unlearning method with the different baseline methods in terms of score (see Section 4.1). The score metric measured the forgetting and utility quality after unlearning. For backdoor verification, the $alpha$ value was multiplied by three up to a maximum of 100%. For actFT to be effective, we multiplied the $alpha$ value for unlearning by the same magnitude. Our method outperformed the previous methods in terms of score with or without BDK for most triggers.
@@ 5 Discussion
We experimented with our algorithm and found specific activation-manipulating formulas that gave us the best unlearning performance for model editing. In Equation 1, negating poisoned activation with BDK and clean activation with $\neg$¬BDK worked the best. With BDK, we negate the activation of the trigger-infected data we want to forget. Previously, Ilharco et al. [22] arrived at the same conclusion as we did, that moving in the negative direction of extracted information can lead to unlearning.
@@ 0.B.3 Effect of Activation Mean and Standard Deviation
Figure 4: Performance of our method when using different values for hyperparameters $\lambda$ and $\gamma$. Compared with or without change in BN MA parameters.